# Improving Diffusion Language Model Decoding through Joint Search in Generation Order and Token Space

### v1

@@ 5.1 Experimental Setup
Baselines.
We compare Order-Token Search (OTS) against several strong baselines: Low-confidence remasking, a greedy decoding method adopted as an optimal base model configuration in Zhao et al. (2025). Random remasking with majority voting, which generates a compute-equivalent set of diverse samples via random remasking and selects answers using a consistency heuristic (Wang et al., 2022). Low-confidence with majority voting, which combines the greedy decoding with the consistency heuristic mentioned above. AR, which follows the left-to-right autoregressive order in generation. AR with majority voting and AR with beam search, which strengthen the AR baseline with, respectively, a consistency heuristic and a search based on OTS likelihood estimator. OTS All Blocks, which scores each expansion by the joint log-likelihood of all blocks revealed so far; OTS Future Blocks, which scores the newly-revealed blocks while masking all unrevealed future blocks (see Appendix A.3). Order Search, a computationally expensive algorithm that uses AR-style likelihood to search for the optimal generation order; Token Search, an equally expensive algorithm that uses AR-style likelihood to search through the top-$K$ likely tokens for each position (see Appendix A.3.3).
@@ 5.2 Order-Token Search Improves Reasoning Accuracy
Diffusion baselines: Order-Token Search improves over remasking and voting on hard reasoning tasks. Within diffusion-style decoding, greedy Low-confidence remasking is already strong, and adding majority voting consistently improves GSM8K averages for both models. Random+MV, which replaces confidence-based remasking with random remasking, can be competitive on GSM8K but substantially degrades tasks requiring more structured reasoning: its Countdown averages drop to 12.7% and 11.4% for LLaDA and LLaDA-1.5, compared with 20.7% and 22.5% for Low-conf+MV. In contrast, Order-Token Search consistently outperforms these diffusion baselines on MATH500, Countdown, and HumanEval. For example, it raises MATH500 averages to 32.8% and 33.8% (vs. 29.7% and 31.3% for Low-conf+MV), improves Countdown to 28.4% and 28.0%, and improves HumanEval to 27.4% and 27.6%.
Autoregressive baselines: Order-Token Search consistently outperforms AR. AR, AR+MV, and AR+beam-search are strong baselines on MATH500, Countdown, and HumanEval, with AR+beam-search achieving the second/third best average performance for both backbones. Nevertheless, Order-Token Search delivers clear gains. Averaging dataset-level accuracies over both LLaDA variants, Order-Token Search attains 71.2/33.3/28.2/27.5% on GSM8K/MATH500/Countdown/HumanEval, compared to 70.0/28.6/13.6/23.5% for AR+MV and 70.0/30.5/19.7/26.1% for AR+beam-search. Notably, AR+beam-search is a $\texttt{block\_size}=1$ special case of our framework, yet it still underperforms full Order-Token Search, underscoring the value of jointly searching over generation orders and token choices rather than performing beam search only in token space.
@@ 5.3 The Necessity of Dedicated Likelihood Estimation
Robust incremental scoring is essential for pruning. Table 1 shows that even with the same OTS expansion and pruning procedure, the choice of likelihood estimator has a large impact. Both OTS All Blocks and OTS Future Blocks underperform Order-Token Search. For LLaDA, Order-Token Search achieves 39.7% overall accuracy, whereas OTS All Blocks and OTS Future Blocks drop to 32.6% and 33.9%, with the largest degradation on Countdown (28.4% $\rightarrow$ 16.2% / 17.4%). The same trend holds for LLaDA-1.5: Order-Token Search reaches 40.4% overall, versus 34.0% / 35.3% for the ablations, again with a large Countdown gap (28.0% $\rightarrow$ 18.4% / 19.2%). Together, these results validate our estimator in Eq. 2: scoring only the newly revealed blocks while conditioning on the model’s full-sequence predictive distribution yields a more stable and discriminative pruning signal than either history-style joint history scoring or scoring without predicted future context.
@@
AR-style likelihood is misaligned with MDMs. Search-based decoding depends on accurate likelihood estimates for pruning, but naive AR-style likelihood is mismatched to MDMs, which are trained to denoise multiple positions in parallel (Section 2.1). In Table 2, Order Search and Token Search instantiate this AR-style estimate by scoring partially denoised sequences via forward passes conditioned only on the currently revealed tokens. Both lag far behind Order-Token Search: on Countdown, Order-Token Search outperforms Order Search by 11%, while Token Search collapses to 0% accuracy. This indicates that directly porting AR beam-search heuristics to MDMs yields unreliable pruning signals; in contrast, Order-Token Search succeeds by pairing joint order–token exploration with a likelihood estimator tailored to incremental denoising.
Figure 4: Countdown accuracy versus test-time compute (NFE) for OTS and majority-voting baselines. For each method, we vary beam size (OTS) or the number of samples (AR+MV, Random+MV), and choose the largest configuration so that all right-most points have roughly matched NFE. At this matched-compute point, OTS with beam size 6 attains 29.3% accuracy, compared to 19.9% for AR+MV and 18.4% for Random+MV, indicating more efficient use of additional FLOPs than simply drawing more independent diffusion samples.
@@ 6 Related Work
Our work addresses this limitation. The iterative denoising of MDMs creates a joint search space over both token values and their generation order, which is inaccessible to autoregressive methods. Our algorithm, Order-Token Search, is designed for this new paradigm, leveraging parallel decoding to explore multiple generation trajectories and select outputs based on overall likelihood.
@@ 7 Conclusion
In this work, we revisited decoding for Diffusion Language Models (DLMs) through the lens of reasoning. We view reasoning as satisfying problem-specific dependencies, which induces a space of valid generation orders; an MDM decoding run corresponds to a single decoding trajectory through this space. Using pass@$k$ as a diagnostic for exploration, we find that standard low-confidence remasking commits to a narrow set of greedy trajectories: it can improve pass@$1$, but yields slow pass@$k$ growth by failing to explore alternative generation orders. Conversely, random remasking broadens trajectory coverage and often increases pass@$k$, but explores blindly and struggles to convert this potential into strong pass@$1$.
To reconcile this trade-off, we introduced Order-Token Search, a decoding algorithm that performs structured search in the joint space of generation orders and token. Order-Token Search maintains multiple partial trajectories, expands them by proposing denoising actions, and prunes using a novel likelihood estimator that scores denoising actions to enable stable selection among competing trajectories. Across mathematical reasoning and coding benchmarks, this joint search yields systematic pass@11gains, matching or surpassing the improvements of expensive post-training methods such as diffu-GRPO. Overall, our results establish joint search over generation order and token content as a key component for unlocking the reasoning capabilities latent in DLMs.
@@ A.1 Order-Token Search Algorithm
1:Input: Prompt $\mathbf{p}$, model $p_{\theta}$, beam size $K$, generation length $L$, total steps $S$, search interval NN, temperature $\tau$, number of blocks $b$.
@@
A.2 Is the Autoregressive Order Fundamentally Superior?
The parallel between random remasking and a fixed AR order in Figure 2 raises a natural hypothesis: perhaps the samples that lead to the high pass@$256$ for random remasking are those that, by chance, follow an autoregressive-like order. If this were true, it would imply that the AR order is a fundamentally superior decoding path for the model.
We test this hypothesis directly with a large-scale correlation study. For a given dataset, we generate 256 samples for each problem using random remasking. For each generated sample, we compute two metrics:
Accuracy: A binary indicator of whether the final solution is correct.
AR Similarity: The Hamming distance between the sequence of positions unmasked in this sample and a canonical left-to-right (AR) order. A low Hamming distance indicates a generation order that is highly similar to AR.
If the hypothesis were correct, we would observe a strong negative correlation between the Hamming distance (difference from AR order) and accuracy; samples that decode in an AR-like order would be more likely to be correct.
Figure S1: Correlation between generation order and accuracy. The x-axis shows how ”chaotic” a generated sample is measured in the Hamming distance of its decoding order from a strict left-to-right (AR) order. The y-axis is the average accuracy of the generated samples with the same chaotic value for a problem. The size of the point represents the number of samples. We find no correlation, indicating that decoding in an AR-like order is not a predictor of success.
Figure S1 plots these results, aggregating data across all samples of the same AR similarity for each problem in a dataset. The result is clear: we find no evidence of a correlation. The coefficient of the fitted lines for each dataset falls under the the order of magnitude of $10^{-3}$. This result holds consistently across all datasets and models we tested.
This null result is profound. It indicates that the autoregressive order is not a uniquely privileged path to a correct solution. Instead, the DLM has learned a rich, multi-faceted solution space where a correct answer can be reached through a vast plurality of different reasoning trajectories. The high pass@$k$ achieved by random remasking is not due to it occasionally stumbling upon an AR order; it is due to the model’s inherent ability to solve problems correctly via many diverse sequences of thought. The AR order’s high pass@$k$ is simply one manifestation of this general capability, not its source.
A.3 Alternative Likelihood Estimators for OTS
@@
A.3.1 OTS All Blocks: Joint Likelihood of All Revealed Blocks
@@
A.3.2 OTS Future Blocks: Likelihood Without Conditioning on Predicted Future
@@
A.3.3 AR-style Likelihood: Left-to-Right Token Log-Likelihood
@@ A.4 Computation Complexity
A.5 Comparison Study of Greedy Decoding and Order-Token Search under Identical Temperature
@@
A.6 Additional Experimental Details
@@
A.6.1 Beam Search Settings
@@
A.6.2 Pass@k Evaluation Settings
@@
A.6.3 Computational Cost Analysis
@@
A.7 Limitations on Sudoku Benchmarks
@@
A.8 Sensitivity of Order-Token Search to Block Size
@@
A.9 Case Study: A Search Instance

### v2

@@ 5.1 Experimental Setup
Baselines. We compare Order-Token Search (OTS) against several strong baselines: Low-confidence remasking, a greedy decoding method adopted as an optimal base model configuration in Zhao et al. (2025). Random remasking, an exploratory decoding method that re-masks uniformly at random. Random remasking with majority voting, which generates a compute-equivalent set of diverse samples via random remasking and selects answers using a consistency heuristic (Wang et al., 2022). Low-confidence with majority voting, which combines the greedy decoding with the consistency heuristic mentioned above. AR, which follows the left-to-right autoregressive order in generation. AR with majority voting and AR with beam search, which strengthen the AR baseline with, respectively, a consistency heuristic and a search based on OTS likelihood estimator. OTS All Blocks, which scores each expansion by the joint log-likelihood of all blocks revealed so far; OTS Future Blocks, which scores the newly-revealed blocks while masking all unrevealed future blocks (see Appendix A.2). Order Search, a computationally expensive algorithm that uses AR-style likelihood to search through the top-$K$ likely positions at every step; Token Search, an equally expensive algorithm that uses AR-style likelihood to search through the top-$K$ likely tokens for each position (see Appendix A.2.3).
@@ 5.2 Order-Token Search Improves Reasoning Accuracy
Diffusion baselines: Order-Token Search improves over remasking and voting on harder tasks. Within diffusion-style decoding, greedy low-confidence remasking is already strong, and adding majority voting consistently improves GSM8K averages for both models. Random+MV, which replaces confidence-based remasking with random remasking, can be competitive on GSM8K but substantially degrades tasks requiring more structured reasoning: its Countdown averages drop to 12.7% and 11.4% for LLaDA and LLaDA-1.5, compared with 20.7% and 22.5% for Low-conf+MV. In contrast, Order-Token Search consistently outperforms these diffusion baselines on harder benchmarks: MATH500, Countdown, and HumanEval. For example, OTS raises MATH500 averages to 32.8% and 33.8% (vs. 29.7% and 31.3% for Low-conf+MV), improves Countdown to 28.4% and 28.0%, and improves HumanEval to 27.4% and 27.6%.
Autoregressive baselines: Order-Token Search consistently outperforms AR. AR, AR+MV, and AR+beam-search are strong baselines on MATH500, Countdown, and HumanEval, with AR+beam-search achieving the second/third best average performance for both backbones. Nevertheless, Order-Token Search delivers clear gains. Averaging dataset-level accuracies over both LLaDA variants, Order-Token Search attains 71.2/33.3/28.2/27.5% on GSM8K/MATH500/Countdown/HumanEval respectively, compared to 70.0/28.6/13.6/23.5% for AR+MV and 70.0/30.5/19.7/26.1% for AR+beam-search. Notably, AR+beam-search is a $\texttt{block\_size}=1$ special case of our framework, yet it still underperforms full Order-Token Search, underscoring the value of jointly searching over generation orders and token choices rather than performing left-to-right beam search only in the token space.
@@ 5.3 The Necessity of Dedicated Likelihood Estimation
Robust incremental scoring is essential for pruning. Table 1 shows that even with the same OTS branching and pruning procedure, the choice of likelihood estimator has a large impact. Both OTS All Blocks and OTS Future Blocks underperform Order-Token Search. For LLaDA, Order-Token Search achieves 39.7% overall accuracy, whereas OTS All Blocks and OTS Future Blocks drop to 32.6% and 33.9%, with the largest degradation on Countdown (28.4% $\rightarrow$ 16.2% / 17.4%). The same trend holds for LLaDA-1.5: Order-Token Search reaches 40.4% overall, versus 34.0% / 35.3% for the ablations, again with a large Countdown gap (28.0% $\rightarrow$ 18.4% / 19.2%). Together, these results validate our estimator in Eq. 2: scoring incremental denoising actions while conditioning on the model’s full-sequence predictive distribution yields a more stable and discriminative pruning signal than either joint history scoring within one forward pass or scoring without predicted future context.
@@
AR-style likelihood is misaligned with MDMs. Search-based decoding depends on accurate likelihood estimates for pruning, but naive AR-style likelihood is mismatched to MDMs, which are trained to denoise multiple positions in parallel (Section 2.1). In Table 2, Order Search and Token Search instantiate this AR-style estimate by scoring partially denoised sequences via forward passes conditioned only on the currently revealed tokens. Both lag far behind Order-Token Search: on Countdown, Order-Token Search outperforms Order Search by 11%, while Token Search collapses to 0% accuracy. This indicates that directly porting AR beam-search heuristics to MDMs yields unreliable pruning signals; in contrast, Order-Token Search succeeds by pairing joint order–token exploration with a likelihood estimation function tailored to incremental denoising.
Figure 4: Countdown accuracy versus test-time compute (NFE) for OTS and majority-voting baselines. For each method, we sweep beam size (OTS) or the number of samples (AR+MV, Random+MV), and set the largest configuration so the right-most points have approximately matched NFE. At this matched-compute point, OTS (beam size 6) achieves 29.3% accuracy, versus 19.9% for AR+MV and 18.4% for Random+MV, demonstrating more efficient returns from additional test-time compute than drawing more independent samples.
@@ 6 Related Work
Our work addresses this limitation. The parallel, order-agnostic decoding of MDMs creates a joint search space over both token values and their generation order, which is inaccessible to autoregressive methods. Our algorithm, Order-Token Search, is designed for this new paradigm, leveraging parallel decoding to explore multiple generation trajectories and select outputs based on overall likelihood.
@@ 7 Conclusion
In this work, we revisited decoding for Diffusion Language Models (DLMs) through the lens of order–token trajectories: the intertwined decisions of where to denoise next and what to write. We identify a clear exploration–exploitation trade-off in standard decoding methods: low-confidence remasking over-exploits local confidence and commits early to a narrow set of greedy trajectories, while random remasking encourages broader trajectory diversity but lacks a reliable selection mechanism to choose good trajectories.
To reconcile this trade-off, we introduced Order-Token Search, which searches in the joint space of generation orders and token values. Order-Token Search maintains partially denoised candidates, branches by proposing diverse denoising actions, and prunes with a likelihood estimator that scores incremental denoising decisions for comparison across competing trajectories. Across mathematical reasoning and coding benchmarks, this joint search yields systematic gains, matching or surpassing improvements typically obtained via post-training methods such as diffu-GRPO. Overall, our results highlight joint order–token search as a key ingredient for converting DLMs’ multi-trajectory capacity into reliable single-sample decoding performance.
@@ A.1 Order-Token Search Algorithm
1:Input: Prompt $\mathbf{p}$, model $p_{\theta}$, beam size $K$, generation length $L$, total steps $S$, search interval $N$, temperature $\tau$, number of blocks $b$.
@@
A.2 Alternative Likelihood Estimators for OTS
@@
A.2.1 OTS All Blocks: Joint Likelihood of All Revealed Blocks
@@
A.2.2 OTS Future Blocks: Likelihood Without Conditioning on Predicted Future
@@
A.2.3 AR-style Likelihood: Left-to-Right Token Log-Likelihood
@@ A.3 Computation Complexity
A.4 Comparison Study of Greedy Decoding and Order-Token Search under Identical Temperature
@@
A.5 Additional Experimental Details
@@
A.5.1 Beam Search Settings
@@
A.5.2 Pass@k Evaluation Settings
@@
A.5.3 Computational Cost Analysis
@@
A.6 Limitations on Sudoku Benchmarks
@@
A.7 Sensitivity of Order-Token Search to Block Size
@@
A.8 Case Study: A Search Instance